OpenStudy (anonymous):

Which ratio correctly compares 60 mm to 45 cm? A. 4 : 3 B. 2 : 15 C. 40 : 3 D. 2 : 150

OpenStudy (anonymous):

@Legends its not that I want the answer I just don't get it :((

OpenStudy (shamim):

Do u know 1cm=10mm

OpenStudy (shamim):

60mm=?cm

OpenStudy (anonymous):

no

OpenStudy (anonymous):

nevermind....

OpenStudy (shamim):

60mm=60/10=?cm

OpenStudy (shamim):

6cm?

OpenStudy (shamim):

6/45=?